OpenStudy (anonymous):

Short Answer. Find the following quotient:x^2-1/x+1

OpenStudy (anonymous):

factor the numerator x^2-1/x+1

OpenStudy (anonymous):

numerator is a perfect square (x-1) (x+1) / x+1

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

\[\huge\frac{x^2-1}{x+1}=\frac{\cancel{(x+1)}(x-1)}{\cancel{(x+1)}}\]

OpenStudy (anonymous):

so what would be the final answer if they all cancel eachother out

OpenStudy (bahrom7893):

they're both the same thing.. the answer's x-1

OpenStudy (anonymous):

They're not all cancelling each other...look above

OpenStudy (anonymous):

only the (x+1)s cancel

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Domain: \[(-\infty, -1)U(-1,\infty)\]

OpenStudy (anonymous):

its just like ( x+1 ) (x - 1) --- -- x ----- = simplify to get your answer ( x+1 ) 1